The length of a rectangular banner is 3 feet longer than its width. If the area is 70 square feet, find the dimensions.
Answer by nerdybill(7384)

The length of a rectangular banner is 3 feet longer than its width. If the area is 70 square feet, find the dimensions.
.
Let w = width of banner
then
w+3 = length of banner
.
w(w+3) = 70
w^2+3w = 70
w^2+3w-70 = 0
factoring:
(w+10)(w-7) = 0
w = {-10, 7}
throw out the negative solution (extraneous) leaving:
w = 7 feet (width of banner)
.
length of banner:
w+3 = 7+3 = 10 feet
.
answer: 7 feet by 10 feet